OpenStudy (anonymous):

What is the volume of a cylinder 20 units high and 10 units in diameter? (a) 78.5 cubic units (b) 314 cubic units (c) 1,571 and 3/7 cubic units (d) 1,570 cubic units ?

OpenStudy (anonymous):

\[V=\pi r^2h\]

OpenStudy (anonymous):

r is radius h is height

OpenStudy (anonymous):

D=2r 10=2r r=5 so: V=(pi)(5^2)20 V=25*20(pi) V=500pi pi=22/7 V=(500)(22/7) V=11000/7 V=1571+3/7

OpenStudy (anonymous):

cubic units

OpenStudy (anonymous):

Thank you! hahaha.

OpenStudy (anonymous):

no problem